OpenStudy (anonymous):

What is the formula for the following geometric sequence? 3, 12, 48, 192, ...

OpenStudy (anonymous):

you always multiply a number by 4 so the nswer is 4x where x is greater than 0

OpenStudy (anonymous):

an = 4 · 3n - 1 an = 3 · (-4)n - 1 an = 3 · 4n - 1 an = (1/4) · 3n - 1 hmmmmmmmmmmmmmmmmmmmmmmm

OpenStudy (anonymous):

this looks strange let me think

OpenStudy (anonymous):

mmmm

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

i not good at secuenses

OpenStudy (anonymous):

sorru

OpenStudy (blurbendy):

the formula for a geometric series is an = a1*r^(n-1) where a1 is the first term in the sequence, and r = common ratio do you know how to find r?

OpenStudy (anonymous):

how

OpenStudy (blurbendy):

r is simply the ratio between the different numbers in the sequence. 3, 12, 48, 192.... You can find r by dividing farthest right number by the one before it 192/48 = 4 48/12 = 4 12/3 = 4 so r = 4

OpenStudy (anonymous):

hmmmm. I understand somewhat. lol thank you though

OpenStudy (blurbendy):

all I did was divide the right term by the one to the left of it starting from 192 Each time, the answer was 4, so that is what r is.

OpenStudy (anonymous):

Got it!

OpenStudy (blurbendy):

cool.

OpenStudy (anonymous):

thanks!

OpenStudy (blurbendy):

any time.